The successor of the predecessor of –50 is (A) –48 (B) –49 (C) –50 (D) –51

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The successor of a whole number is the number obtained by adding 1 to it.
To find the predecessor of a number, subtract one from the given number.
Thus, predecessor of–50 = - 50–1 = - 51
Then, successor of–51 = - 51 + 1 = 50
Thanks